OpenStudy (anonymous):

find the zeros of the function y=x^2-4x+3 by factoring

geerky42 (geerky42):

x²-4x+3 = (x - ?)(x - ?)

OpenStudy (anonymous):

(x-3)(x-1)

geerky42 (geerky42):

Excellent. Now find zero. Do you know how to?

OpenStudy (anonymous):

i have no idea

geerky42 (geerky42):

(x-3)(x-1) = 0 so either (x - 3) or (x - 1) must be equal zero.

OpenStudy (anonymous):

(x-3)

geerky42 (geerky42):

or (x - 1). So set both terms equal to zero. (x - 3) = 0 (x - 1) = 0 Find x. You will have two answers.

OpenStudy (anonymous):

x=3 x=1

OpenStudy (anonymous):

thank you